Hehe:

For what values of θ, 0° ≤θ≤ 360°, does y=cos(θ) equal 0?

TheSmartOne:

Hello there! Do you know how the graph of y=cos\( (\theta )\) looks like? If not, here you go: https://www.desmos.com/calculator/b2gf1c4m1a 360 degrees is the same thing as \( 2\pi\) radians, if you didn't know that either. Do you see which values of \(\theta\) give you 0? You can convert from radians to degree by multiply it with the following conversion unit \(\Large \frac{360^o}{2\pi}\)